Solve the system using substitution:
3x=2y-6(1)/(2)
4x+y=6

Answer by lwsshak3(11628)

Solve the system using substitution:
3x=2y-6(1)/(2)
4x+y=6
y=6-4x
3x=2(6-4x)-3
3x=12-8x-3
11x=9
x=9/11